Wersja testu A 19 czerwca 2002r.

1. a. N b. N c. T d. N

2. a. N b. T c. N d. N

3. a. N b. T c. T d. N

4. a. N b. N c. T d. T

5. a. N b. N c. T d. T

6. a. T b. N c. T d. T

7. a. T b. T c. N d. N

8. a. T b. N c. T d. N

9. a. T b. T c. N d. N

10. a. T b. T c. T d. N

11. a. N b. T c. N d. N

12. a. T b. N c. N d. N

13. a. T b. T c. T d. T

14. a. N b. N c. T d. N

15. a. N b. N c. N d. N

16. a. N b. N c. T d. N

17. a. N b. N c. T d. T

18. a. T b. N c. N d. N

19. a. N b. N c. N d. T

20. a. T b. T c. N d. wycofane